HIGH RESOLUTION FOURIER TRANSFORM UV EMISSION SPECTROSCOPY OF THE 410 NM BAND OF THE TiCl RADICAL

\underline{TAKASHI IMAJO}, DONBBING WANG, KEIICHI TANAKA, and TAKEHIKO TANAKA, Department of Chemistry, Faculty of Science, Kyushu University 33, Hakozaki, Higashi ku, Fukuoka 812-8581, Japan.

High resolution UV emission spectrum of the 410 nm band of the TiCl radical was measured by Fourier transform spectrometer (Bruker IFS120HR) with the resolution of 0.04 cm^{−1}. This band was assigned as the electronic transition of ^{2}Φ −^{2} Δ band of the TiCl radical by Phillips and Davis [1], but recent study of ZrCl[2] suggested that this band might be ^{4}Γ − X^{4}Φ or ^{4}Φ − X^{4}Φ band. In the present work a new electronic assignment of ^{4}Γ - X ^{4}Φ was proposed. The vibrational assignments were made for each subbands. The vibrational bands of (0-0), (1-1), (2-2), (3-3), and (4-4) were assigned for the subbands of 4Γ_{5/2} −^{4} Φ_{3/2} and ^{4}Γ_{7/2} −^{4} Φ_{5/2}. For the subband of ^{4}Γ_{9/2} −^{4} Φ_{7/2}, the (0-0), (1-1) and (2-2) bands were assigned, but for the ^{4}Γ_{11/2} −^{4} Φ_{9/2} subband, only (0-0) band was assigned. Rotational analysis were done for ^{4}Γ_{5/2} −^{4} Φ_{3/2} and 4Γ_{7/2} −^{4} Φ_{5/2} subbands. The band origins, rotational constants, and centrifugal distortion constants were determined by the least square method. The rotational constants were consistent with those reported by Ram and Bernath [3] within the experimental error. ¿From the line intensities of P- and R- branches, ΔΛ was derived to be +1, and therefore upper state was determined to be ^{4}Γ. Further measurements and analysis for the subbands of ^{4}Γ_{9/2} −^{4} Φ_{7/2} and ^{4}Γ_{11/2} −^{4} Φ_{9/2} should be needed to confirm the present electronic assignment definitely.

1. J. G. Phillips and S. P. Davis, Astr. Phys. J. Suppl. Ser 71, 163-172 (1989).

2. R. S. Ram and P. F. Bernath, J. Mol. Spectrosc. 186, 335-348 (1997).

3. R. S. Ram and P. F. Bernath, J. Mol. Spectrosc. 186, 113-130 (1997).